Examine whether the second order partial derivatives of f at (0,0) exist or not if


f :R² ➡R is defined by


f(x, y) ={x²y/√x+y² , xy≠0 and 0, xy=0


1
Expert's answer
2022-01-25T16:02:06-0500

"\text{By definition }f:\R\rightarrow\R,\text{defined by,}\\
f(x,y)=
\begin{cases} 
\frac{x^2y}{\sqrt{x+y^2}} & \text{if}\ (x,y)\neq(0,0)\\0 & \text{if\ (x,y)}=(0,0) 
\end{cases}\\
\text{Now,}\\
f(x,0)=f(0,y)=0\\
\text{Thus, }f_x(x,0)=f_y(0,y)=0\\
\Rightarrow f_{xx}(x,0)=f_{yy}(0,y)=0\\
\text{Thus, the second partial derivatives of the given function exists at (0,0) and are: }\\
f_{xx}(0,0)=f_{yy}(0,0)=f_{xy}(0,0)=0"